# Mutual Enhancement of Large Language and Reinforcement Learning Models through Bi-Directional Feedback Mechanisms: A Case Study

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities for reinforcement learning (RL) models, such as planning and reasoning capabilities. However, the problems of LLMs and RL model collaboration still need to be solved. In this study, we employ a teacher-student learning framework to tackle these problems, specifically by offering feedback for LLMs using RL models and providing high-level information for RL models with LLMs in a cooperative multi-agent setting. Within this framework, the LLM acts as a teacher, while the RL model acts as a student. The two agents cooperatively assist each other through a process of recursive help, such as ”I help you help I help.” The LLM agent supplies abstract information to the RL agent, enabling efficient exploration and policy improvement. In turn, the RL agent offers feedback to the LLM agent, providing valuable, real-time information that helps generate more useful tokens. This bi-directional feedback loop promotes optimization, exploration, and mutual improvement for both agents, enabling them to accomplish increasingly challenging tasks. Remarkably, we propose a practical algorithm to address the problem and conduct empirical experiments to evaluate the effectiveness of our method.
@@ 1 Introduction
Large Language Models (LLMs) openai2023gpt4; chang2023survey have showcased exceptional prowess across various domains. Notably, LLMs find applications in furnishing information for tasks like robot planning singh2023progprompt, machine translation zhang2023prompting and medicine thirunavukarasu2023large. In parallel, RL has demonstrated remarkable capabilities in various domains, including achieving human-level performance in games such as the game of Go silver2016mastering and multiplayer poker brown2019superhuman. LLMs have been increasingly incorporated to enhance the performance of Reinforcement Learning (RL) du2023guiding; szot2023large. Likewise, RL has also been employed to augment the capabilities of LLMs, furthering their effectiveness ouyang2022training. Nevertheless, the effective harnessing of LLMs’ latent potential in solving complex tasks, through the synergistic integration with powerful RL frameworks sutton2018reinforcement, remains a formidable challenge.
The research most related to our study includes the works of Carta et al. carta2023grounding and Tran et al. tran2023exploring. In the work of Carta et al. carta2023grounding, they employ LLMs as RL policies to acquire task-solving capabilities while learning new knowledge through interactive experiences. Their experimental findings suggest that their method outperforms baseline approaches. However, a potential limitation of their work is the absence of instruction feedback from RL models, which may impact the overall effectiveness of their method. In the work of Tran et al. tran2023exploring, they deploy RL to train a conversational agent using a simulator and an initial text generated by a generative chat agent. Subsequently, they input the data from the RL-trained agent to the generative chat agent. Although their experiment results indicate that their method performs better than baselines, a concern of this approach is the potential time consumption associated with RL training for multi-turn conversations, as each conversation may necessitate RL training requests. Additionally, achieving self-online learning for task execution could be challenging in this framework.
@@
In this study, to address the above challenge, we propose a teacher-student learning framework in a cooperative game, where the integration of RL models (students) and LLMs (teachers) with bi-directional feedback gu2023human may be an effective solution. The two models cooperatively to carry out complex tasks, which can be considered a win-win collaboration, where the RL model and the LLM act as two agents, cooperating to complement, assist, and provide feedback to each other, ultimately solving the problem together.
@@ 2 Method
LLMs (teachers) help RL models (students): While it is often challenging for LLMs to provide instructions encompassing perfect and comprehensive environmental information, LLMs can supply RL models with approximated information. Providing such rudimentary guidance by LLMs serves the purpose of streamlining the exploration process undertaken by RL models. Consequently, this streamlined exploration process yields a discernible reduction in the exploration space and the time required for RL models to ascertain and establish an optimal policy. This phenomenon underscores the potential utility of LLMs in mitigating the challenges associated with imperfect instructional input, thereby contributing to the enhanced efficiency of RL models in their quest to identify optimal policies.
RL models (students) help LLMs (teachers): During the execution of a policy within the RL framework, RL models benefit from the support provided by LLMs. In this collaborative process, RL models are not only recipients but also evaluators of the output generated by LLMs. This reciprocal interaction allows RL models to offer constructive feedback to LLMs, thereby facilitating an iterative refinement of LLMs’ performance. LLMs progressively acquire a more nuanced understanding of the underlying environments with the progression of iterations. Consequently, they become increasingly adept at furnishing improved output, which aids RL models and LLMs in executing complex tasks with greater efficacy. This iterative and symbiotic relationship between RL models and LLMs emphasizes the potential for continuous improvement and optimization in their collaborative endeavors. The corresponding practical algorithm is provided in Algorithm 1.
@@ 3 Experiments
Our experimental investigation is conducted within the context of the BabyAI benchmark babyaiiclr19. To facilitate our experimentation, we leverage the Lamorel framework carta2023grounding. Within the Lamorel framework, we augment LLMs with RL instruction feedback, thereby establishing a feedback loop for the enhancement of LLM performance. Specifically, we proceed to conduct experiments focusing on the task of GoToRedBallNoDists-v0. The experiments are executed over two cases: one comprising 40 iteration steps and another spanning 2100 iteration steps. In the evaluation, we draw a comparative analysis between our proposed method and the baseline. In our method, we consider RL model feedback to LLMs and LLMs’s information to RL models. However, the state-of-the-art baseline represented by the original Lamorel method that lacks such instruction feedback to LLMs from RL models.
It is noteworthy that for our experiments, we employ the ”google/flan-t5-small” model chung2022scaling as the LLM, characterized by a parameter count of 80 million. The experimental results are presented in Figure 2. These findings clearly illustrate the superior performance of our method, as quantified by the performance value metric (where higher values indicate better performance). Furthermore, our method demonstrates notably expedited convergence (one-shot/few-shot learning) when compared to the Lamorel baseline. This empirical evidence highlights the effectiveness of our approach in harnessing bi-directional feedback between RL models and LLMs for improving performance in the context of the BabyAI benchmark.
Figure 2: Experiments on BabyAI tasks babyaiiclr19 with 40 (a) and 2100 (b) iteration steps.
@@ 4 Conclusion
In this study, we developed a teacher-student learning framework for unlocking LLMs’ powerful capabilities by leveraging an RL model with bi-directional feedback mechanisms in a cooperative game setting. To empirically assess the effectiveness of our method, we conducted experiments using the BabyAI benchmark as an assessment platform. The results of these experiments demonstrate the superior performance of our approach in comparison to the state-of-the-art baseline, underscoring its potential for substantially enhancing learning outcomes. Importantly, our approach could exhibit promise in cultivating safe and robust learning systems gu2022review, particularly when confronted with the inherent challenges of imperfect information environments. Furthermore, we hope our insights inspire novel avenues of research in the realms of LLMs and RL for future investigations.

### v2

# Mutual Enhancement of Large Language and Reinforcement Learning Models through Bi-Directional Feedback Mechanisms: A Planning Case Study

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities for reinforcement learning (RL) models, such as planning and reasoning capabilities. However, the problems of LLMs and RL model collaboration still need to be solved. In this study, we employ a teacher-student learning framework to tackle these problems, specifically by offering feedback for LLMs using RL models and providing high-level information for RL models with LLMs in a cooperative multi-agent setting. Within this framework, the LLM acts as a teacher, while the RL model acts as a student. The two agents cooperatively assist each other through a process of recursive help, such as “I help you help I help.” The LLM agent supplies abstract information to the RL agent, enabling efficient exploration and policy improvement. In turn, the RL agent offers feedback to the LLM agent, providing valuable, real-time information that helps generate more useful tokens. This bi-directional feedback loop promotes optimization, exploration, and mutual improvement for both agents, enabling them to accomplish increasingly challenging tasks. Remarkably, we propose a practical algorithm to address the problem and conduct empirical experiments to evaluate the effectiveness of our method.
@@ 1 Introduction
Large Language Models (LLMs) [13, 4] have shown exceptional performance across various domains. Notably, LLMs are useful for applications like robot planning [17], machine translation [24] and medicine [20]. In parallel, RL has demonstrated remarkable capabilities in various domains, including achieving human-level performance in games such as the game of Go [16] and multi-player poker [2]. LLMs have been increasingly incorporated to enhance the performance of RL [8, 19]. Likewise, RL has also been employed to augment the capabilities of LLMs, furthering their effectiveness [14, 9]. Nevertheless, the effective harnessing of LLMs’ latent potential in solving complex tasks, through the synergistic integration with powerful RL frameworks [18], remains a formidable challenge. Therefore, a critical question that emerges in this field is: How does an RL model cooperate with LLMs to perform a given task effectively?
Facilitating cooperation between RL models and LLMs requires mutually beneficial actions, leading to enhancing each model’s performance. However, it is challenging to meet these needs due to RL models’ and LLMs’ distinct decision-making characteristics. Specifically, the capabilities of LLMs generally exceed those of RL models, indicating the importance of developing methodologies for instructing RL models to acquire high-level knowledge and ensuring RL models can deliver real-time feedback to LLMs.
@@
In this study, to address the above challenge, we propose a teacher-student learning framework in a cooperative game, where the integration of RL models (students) and LLMs (teachers) with bi-directional feedback [10] may be an effective solution. The two models cooperatively carry out complex tasks, which can be considered a win-win collaboration, where the RL model and the LLM act as two agents, cooperating to complement, assist, and provide feedback to each other, ultimately solving the problem together.
2 Related Work
The relation between RL models and language representation is investigated in several methods [1, 5, 9, 12, 14, 15, 22, 23, 25]. For instance, in the work of [chen2023introspective], they deploy LLMs as an optimization objective in an RL decision-making loop. Then, they further make LLMs reflect their decision results to refine LLMs’ output using prompt engineering tips. In the work of [yang2021safe], they leverage text as safety constraints for RL safe exploration. In the work of [ouyang2022training], they train LLMs to align with human values in an RL process, where the reward model is a supervised label to teach LLMs to follow human instructions.
The research most related to our study includes the works of [carta2023grounding] and [tran2023exploring]. In the work of [carta2023grounding], they employ LLMs as RL policies to acquire task-solving capabilities while learning new knowledge through interactive experiences. Their experimental findings suggest that their method outperforms baseline approaches. However, a potential limitation of their work is the absence of instruction feedback from RL models, which may impact the overall effectiveness of their method. In the work of [tran2023exploring], they deploy RL to train a conversational agent using a simulator and an initial text generated by a generative chat agent. Subsequently, they input the data from the RL-trained agent to the generative chat agent. Although their experiment results demonstrate that their method performs better than baselines, a concern of this approach may be the potential time consumption associated with RL training for multi-turn conversations, as each conversation may necessitate RL training requests. Additionally, achieving self-online learning for task execution could be challenging in this framework.
@@ 3 Method
LLMs (teachers) help RL models (students): LLMs often struggle to generate instructions that fully capture detailed and precise environmental information. However, they can still provide approximate guidance to RL models, aiding the exploration process. By narrowing the exploration space and accelerating policy discovery, such guidance improves the efficiency of RL training. This highlights the potential of LLMs to mitigate challenges arising from imperfect instructions, thereby enhancing RL performance in policy optimization.
RL models (students) help LLMs (teachers): During policy execution in the RL framework, RL models benefit from the guidance provided by LLMs. In this collaborative process, RL models not only utilize but also evaluate the outputs generated by LLMs. This reciprocal interaction enables RL models to provide constructive feedback, facilitating the iterative refinement of LLM performance. As iterations progress, LLMs can better understand the environment, allowing them to generate increasingly effective guidance. This, in turn, enhances the ability of both RL models and LLMs to tackle complex tasks with greater efficiency. The iterative relationship between RL models and LLMs highlights their potential for continuous improvement and optimization. The corresponding algorithm is presented in Algorithm 1.
@@ 4 Experiments
Our experimental study is conducted within the BabyAI benchmark [babyaiiclr19], utilizing the Lamorel framework [carta2023grounding] to support our investigation. Within this framework, we integrate RL instruction feedback into LLMs, establishing a feedback loop to enhance LLM performance. Specifically, we focus on the GoToRedBallNoDists-v0 planning task, conducting experiments under two conditions: one with 40 iteration steps and another with 2100 iteration steps. We perform a comparative analysis between our proposed method and the baseline for evaluation. Our approach incorporates bidirectional interaction, where RL models provide feedback to LLMs, and LLMs supply information to RL models. In contrast, the state-of-the-art baseline, represented by the original Lamorel method, lacks this feedback mechanism from RL models to LLMs.
It is noteworthy that for our experiments, we employ the ”google/flan-t5-small” model [chung2022scaling] as the LLM, characterized by a parameter count of 80 million. The experimental results are presented in Figure 2. These findings clearly illustrate the superior performance of our method, as quantified by the performance value metric (where higher values indicate better performance). Furthermore, our method demonstrates notably expedited convergence (one-shot/few-shot learning) when compared to the Lamorel baseline. This empirical evidence highlights the effectiveness of our approach in harnessing bi-directional feedback between RL models and LLMs to improve performance in the context of the BabyAI benchmark.
Figure 2: Experiments on BabyAI tasks [babyaiiclr19] with 40 (a) and 2100 (b) iteration steps.
@@ 5 Conclusion
In this study, we developed a teacher-student learning framework for unlocking LLMs’ powerful capabilities by leveraging an RL model with bi-directional feedback mechanisms in a cooperative game setting. To empirically assess the effectiveness of our method, we conducted experiments using the BabyAI benchmark as an assessment platform. The results of these experiments demonstrate the superior performance of our approach in comparison to the state-of-the-art baseline, highlighting its potential for substantially enhancing learning outcomes. Notably, our approach holds promise for fostering safe and robust learning systems [gu2022review], particularly in environments characterized by imperfect information. Furthermore, we hope that our findings inspire novel research directions at the intersection of LLMs and RL. As part of future work, we plan to extend our method to more challenging tasks and assess its effectiveness in complex real-world applications.